OpenStudy (anonymous):

Complete conversation with the right possession Tu: Ah, Carlos es ---- hermanastro. No lo sabia ?Y la chica esta a su lado es tu hermana? Raquel: No. Carlos esta castado y esa chica es --- esposa, Carmen. ---- hermana esta sentada con ---- amigos. Tu: ?Y los senores que estan detras son ----pares, Raquel? Raquel: No. No son ---- padres. Son los padres de Carmen. Los ---- estan sentados ahi delante. Tu: !Ah! Entonces esa es la familia politica de Carlos. Raquel: Si, si, esos senores son ---- suegros y la chica del pelo largo es ---- cunada.

OpenStudy (anonymous):

can somebody please help me my teacher went over this the last five minutes of class and I'm confuse

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Possessive adjectives: Yo: mi = I: my Tú: tu = You: your Él/ella/usted: su = He:his/She:her (In this case, usted=you and su= your. In some countries, "usted" is more formal than "tú". In Spanish "It" as a pronoun doesn't exist, but "its" can be translated to "su" too). Nosotros: nuestro(a)(os)(as)= possible endings / We:our Vosotros: vuestro(a)(os)(as)= possible endings / You:your (plural) Ellos/ellas/ustedes: sus / They: their (In this case, "ustedes" corresponds to "you" in the plural form, then "sus"=your). Possessive pronouns: Yo: mío/a/os/as (possible endings) = I: mine Tú: tuyo/a/os/as (possible endings) Él/ella/usted: suyo/a/os/as = He:his/she:hers/you:yours (same case mentioned)/it:its Nosotros: nuestro/a/os/as = We: ours Vosotros: vuestro/a/os/as = You: yours (plural) Ellos/ellas/ustedes: suyo/a/os/as = They: their [Ustedes:you (plural)] Well, now don't forget to write a tilde when a word requires it, because, apart of to be a grammar rule, this can change the meaning of the word: Tú (personal pronoun), tu (possessive adjective), sabía (saber in past tense), sabia (wise person), esta (demonstrative adjective), está (verb estar conjugated for él, ella, usted in present tense), detras (doesn't exist), detrás (behind), ahi (doesn't exist), ahí (there), politica (a word peculiar of Costa Rica to refer to a ridiculous woman or man= politico), política (policy/in-laws), si (if), sí (yes). And remember the "Ñ, ñ" is a very important letter is the Spanish alphabet, then don't omit the mark over the "n": cunada (doesn't exist), cuñada (sister-in-law), senores/senor (doesn't exist), señores/señor (mister/s or sir/s). Other corrections: the inverted question mark (at the beginning of a sentence) is this: "¿", and the inverted exclamation mark is this: "¡". And be carefull with the spelling: pares (pairs), padres (parents). OK, when you have a noun after the 'possessive word' you need, then you have to write a possessive adjective, if there isn't a noun, you have to substitute it with a possessive pronoun. Maybe these sites can help you too: http://www.studyspanish.com/lessons/possadj.htm http://www.studyspanish.com/lessons/posspro.htm Go ahead! If you need more help, let me know.